OpenStudy (anonymous):

Find the Integral Integral of (pi/2 , 3pi/2) Cosx/Sin^2x dx

hartnn (hartnn):

Hi :) did you try u = sin x ? du =...?

OpenStudy (anonymous):

let me try that yes, I totally forgot I had that option. Um, u= sin^2x ? Or should I omit the ^2? and the du= cosxdx?

hartnn (hartnn):

yessss u = sin x du = cos x dx is correct now change the limits too when x =pi/2 u=...? when x= 3pi/2 u=...?

OpenStudy (anonymous):

x= pi/2 -> u= sin(pi/2) ? and then the same with 3pi/2?

hartnn (hartnn):

yes, whats u = sin(pi/2) = ...? whats u = sin(3pi/2) = ...?

OpenStudy (anonymous):

1 and -1

hartnn (hartnn):

yes, so your new limits are from 1 to -1

hartnn (hartnn):

whats your new integral ?

OpenStudy (anonymous):

integral of (-1, 1) u^2 / du ?

hartnn (hartnn):

u = sin x and sin x is in denominator , so \(\Large \int \limits_{-1}^1\dfrac{1}{u^2}du\) right ??

OpenStudy (anonymous):

yes, then substitute with the u values, and then substitute with the integral values ?

hartnn (hartnn):

integrate 1/u^2 can you ?

OpenStudy (anonymous):

1/ sinx^2 so then it would be sin(-1)^2

hartnn (hartnn):

there is no sin x now...

hartnn (hartnn):

your new integral is just in u

OpenStudy (anonymous):

how is that? u= -1?

hartnn (hartnn):

 here n = -2